OpenStudy (anonymous):

Factor the binomial completely x^2-1/4

OpenStudy (skullpatrol):

any ideas?

OpenStudy (skullpatrol):

$$\Huge a^2-b^2$$

OpenStudy (skullpatrol):

$$\Huge a^2-b^2=(a+b)(a-b)$$

OpenStudy (skullpatrol):

$$\Huge x^2-\dfrac{1}{4}=x^2-(\dfrac{1}{2})^2$$

OpenStudy (anonymous):

(x+1)(x-1)

OpenStudy (skullpatrol):

try again @MIS08

OpenStudy (skullpatrol):

b = 1/2